The following fractions represent just three different numbers. Separate them into three groups of equivalent fractions, by changing each one to its simplest form.
3/18

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3 / 18 = (1 × 3) / (6 × 3)
=1 / 6

​Thanks